# Effective Polaron Dynamics of an Impurity Particle Interacting with a Fermi Gas

### v1

@@ 2.3 Almost-bosonic operators and patch decomposition
Similarly to (2.11) introduced in the previous subsection, we define
$c^{\ast}(\eta)\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\eta_{\alpha}(k)}c_{\alpha}^{\ast}(k)$ (2.20)
with inner product $\langle\eta,\varphi\rangle\coloneqq\sum_{k\in\Gamma^{\textnormal{nor}}}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\eta_{\alpha}(k)}\varphi_{\alpha}(k)$ for all $\eta,\varphi\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$.
@@
  • The Canonical Anticommutation Relations (CAR) are satisfied up to an error term (see [BNP+19, Lemma 4.1]): It holds for all $k^{\prime},k\in\Gamma^{\textnormal{nor}}$ and $\alpha\in\mathcal{I}_{k},\beta\in\mathcal{I}_{k^{\prime}}$
@@
Also note that the summation in the definition of the almost-bosonic operators $c_{\alpha}^{\ast}(k)$ and $c_{\alpha}(k)$ involves only finite sets. Unlike in the exactly bosonic case our almost-bosonic operators therefore inherit boundedness from the fermionic constituents which satisfy $\|a_{k}^{\ast}\|=\|a_{k}\|=1$. Subtle questions about the self-adjointness and domain of the almost-bosonic operators remain trivial in our case.
2.4 Almost-bosonic coherent state
Since $B\coloneqq c^{\ast}(\eta)-c(\eta)$ defines for all $\eta\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ a bounded operator and satisfies $B=-B^{\ast}$, the exponential operator $e^{B}$ is well-defined and is unitary.
@@ 2.4 Almost-bosonic coherent state
Define for $\eta\in\bigoplus_{k\in\Gamma^{\textnormal{nor}}}l^{2}(\mathcal{I}_{k})$ the Weyl operator
@@
i.e. the coupled coherent state corresponds to a superposition of different particle number. Later on , the terms $c^{\ast}(\eta)$ and $c(\varphi)$ with the function $(\eta)_{\alpha}(k)\equiv(\eta^{y})_{\alpha}(k)\equiv e^{iky}\varphi_{\alpha}(k)$ will take the role of interaction term with an multiplication operator acting on $\phi$. Thus the occurrence of $\phi$ in each component of the vector representation stresses the coupling between the $L^{2}$ function $\phi$ and the bosonic $n$-particle state via the interaction $\eta$.

### v2

@@ 2 Preliminaries
Similarly to (2.11) introduced in the previous subsection, we define as in [BNP+21a]
$c^{\ast}(\eta)\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\eta_{\alpha}(k)}c_{\alpha}^{\ast}(k)$ (2.20)
with inner product $\langle\eta,\varphi\rangle\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\overline{\eta_{\alpha}(k)}\varphi_{\alpha}(k)$ for all $\eta,\varphi\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$.
@@
  • The Canonical Commutation Relations (CCR) are satisfied up to an error term (see [BNP+19, Lemma 4.1]): It holds for all $k^{\prime},k\in\Gamma$ and $\alpha\in\mathcal{I}_{k},\beta\in\mathcal{I}_{k^{\prime}}$
@@
Also note that the summation in the definition of the almost-bosonic operators $c_{\alpha}^{\ast}(k)$ and $c_{\alpha}(k)$ involves only finite sets. Unlike in the exactly bosonic case our almost-bosonic operators therefore inherit boundedness from the fermionic constituents which satisfy $\|a_{k}^{\ast}\|=\|a_{k}\|=1$. Subtle questions about the mutual adjointness and domain of the almost-bosonic operators remain trivial in our case.
3 Main results
3.1 Effective time evolution
$\mathbb{H}^{\text{eff}}\coloneqq$ $-\beta\Delta_{y}+\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)c_{\alpha}(k)+E_{N}^{\text{pw}}$
$+\lambda\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\hat{V}(k)e^{iky}n_{\alpha}(k)\bigg{(}c_{\alpha}^{\ast}(k)+c_{\alpha}(-k)\bigg{)}$ (3.1)
be our effective Hamiltonian with $\epsilon_{\alpha}(k)=2k_{F}|k\cdot\omega_{\alpha}|$ and $E_{N}^{\text{pw}}$ as defined in (2.5). We introduce for all $k\in\Gamma$ and $\alpha\in\mathcal{I}_{k}$
with $\Phi(h_{y})\coloneqq c^{\ast}(h_{y})+c(h_{y})$ and
$c^{\ast}(h_{y})$ $\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}(h_{y})_{\alpha}(k)\ c_{\alpha}^{\ast}(k),$ (3.4)
$c(h_{y})$ $\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\overline{(h_{y})_{\alpha}(k)}\ c_{\alpha}(k),$ (3.5)
$\mathbb{D}_{\text{B}}$ $\coloneqq\sum_{k\in\Gamma}\sum_{\alpha\in\mathcal{I}_{k}}\epsilon_{\alpha}(k)c_{\alpha}^{\ast}(k)c_{\alpha}(k)\quad\text{with }\epsilon_{\alpha}(k)=2k_{F}|k\cdot\omega_{\alpha}|$ (3.6)
To state an effective description of the time evolution of those excitations we compare the particle-hole transformed microscopic dynamics $R^{\ast}e^{-i\mathbb{H}t}R\psi$ with the effective time evolution $e^{-i\mathbb{H}^{\textnormal{eff}}t}\psi$ in Hilbert space norm:
Theorem 3.1 (Effective dynamics of the system).
Assume that $\hat{V}\geq 0$ is compactly supported and satisfies $\hat{V}(-k)=\hat{V}(k)$ for all $k\in\mathbb{Z}^{3}$. Let $\lambda\in[k_{\textnormal{F}}^{-1/6},1]$ be the interaction parameter as introduced in (1.2) and take the number of patches to be $M=N^{\frac{16}{45}}$ with $\delta=\frac{2}{15}$ as introduced in (2.15). Then it holds for the initial state $\psi=\phi\otimes\Omega\in L^{2}(\Lambda,\textnormal{d}y)\otimes\mathcal{H}_{N}^{-}$ that there is a $C>0$ depending only on the interaction $V$ such that for all $k_{\textnormal{F}}\geq 2$ and $t\geq 0$
$\|R^{\ast}e^{-i\mathbb{H}t}R\psi-e^{-i\mathbb{H}^{\textnormal{eff}}t}\psi\|\leq C\left(e^{C\lambda k_{\textnormal{F}}t}-1\right)k_{\textnormal{F}}^{-\frac{1}{5}}.$
Remark 3.2.
Note that the right hand side of the bound is indeed small as long as $t\in\mathcal{O}(k_{\textnormal{F}}^{-1}\lambda^{-1})$. The error $k_{\textnormal{F}}^{-1/5}$ is a result of the optimized choice of $M$ and $\delta$. The non-optimal error bound depends on the patch parameters and is given by
$C\left((1+\lambda^{-1})\lambda^{-1}(M^{-\frac{1}{2}}+MN^{-\frac{2}{3}+\delta})+(N^{-\frac{\delta}{2}}+N^{-\frac{1}{6}}M^{\frac{1}{4}})\right)\left(e^{C\lambda k_{\textnormal{F}}t}-1\right).\$ (3.7)
3.2 Effective coherent state
The effective time evolved state from Theorem 3.1 can be further simplified. Our second result shows how the dynamics can be approximated on the level of states. In order to state the second main result, we introduce almost-bosonic coherent states.
Note that since $B\coloneqq c^{\ast}(\eta)-c(\eta)$ defines for all $\eta\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ a bounded operator and satisfies $B=-B^{\ast}$, the exponential operator $e^{B}$ is well-defined and is unitary.
@@ 3.2 Effective coherent state
Define for $\eta\in\bigoplus_{k\in\Gamma}l^{2}(\mathcal{I}_{k})$ the Weyl operator
@@
i.e. the coupled coherent state corresponds to a superposition of different particle number.
Remark 3.5.